# On Sum-Free Functions

### v1

@@ 4. Generalizations
Recall that a function $f:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is said to be $k$th order sum-free if $\sum_{x\in A}f(x)\neq 0$ for all $k$-dimensional $\mathbb{F}_{q}$-affine subspaces $A$ of $\mathbb{F}_{q^{n}}$. In general, compared with their binary counterpart, $q$-ary sum-free functions are more difficult to characterize, construct and analyze. For example, there does not seem to be a $q$-ary version of APN functions that represents the 2nd order $q$-ary sum-free functions. We first take a look of the 1st order and the $(n-1)$th order $q$-ary sum-free functions.
@@ 4.1. The 1st order
Let $f(X)=\sum_{i=0}^{q^{n}-1}a_{i}X^{i}\in\mathbb{F}_{q^{n}}[X]$. Then $f:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is $1$st order sum-free if and only if
@@
4.2. The $(n-1)$th order
@@
$\kern 85.35826pt\cdot\sum_{i=0}^{q^{n}-1}a_{i}\sum_{x\in\mathbb{F}_{q^{n}}}x^{i+j_{0}q^{0}+\cdots+j_{n-1}q^{n-1}}$
@@
In the above notation, $f$ is $(n-1)$th order sum-free if and only if
@@
Condition (4.5) can always be satisfied since the right side of (4.5) takes at most $q^{n}-1$ values as $b$ runs over $\mathbb{F}_{q^{n}}^{*}$. Condition (4.6) is difficult to verify in general. However, in special cases, (4.6) can be made more explicit.
Let $10<t\leq q-$ 1 and assume that for $10<j_{0}+\cdots+j_{n-1}\leq q-$ 1,
@@
$\binom{j_{0}+\cdots+j_{n-1}}{j_{0},\dots,j_{n-1}}=0,\ \text{i.e., the sum $j_{0}+\cdots+j_{n-1}$ has no carries in base $p$,}$
@@
[…]
Then the left side of (4.6) becomes
@@
is a $q$-polynomial. Therefore, if $L_{t}(X)$ is a permutation polynomial of $\mathbb{F}_{q^{n}}$, then (4.6) is satisfied.
@@
4.3. The Moore determinant and related formulas
Let $X_{1},\dots,X_{k}$ be independent indeterminates, and define
$\Delta(X_{1},\dots,X_{k})=\left|\begin{matrix}X_{1}&\cdots&X_{k}\cr X_{1}^{q}&\cdots&X_{k}^{q}\cr\vdots&&\vdots\cr X_{1}^{q^{k-1}}&\cdots&X_{k}^{q^{k-1}}\end{matrix}\right|\in\mathbb{F}_{q}[X_{1},\dots,X_{k}],$
which is known as the Moore determinant over $\mathbb{F}_{q}$.
Lemma 4.3 ([13, Lemma 3.51]).
We have
Corollary 4.4.
$v_{1},\dots,v_{k}\in\mathbb{F}_{q^{n}}$ are linearly independent over $\mathbb{F}_{q}$ if and only if $\Delta(v_{1},\dots,v_{k})\neq 0$.
For $0\leq i\leq k$, let
@@ 4.3. The Moore determinant and related formulas
(i) $\Delta(Y,X_{1},\dots,X_{k})$ is a polynomial of degree $q^{k}$ in $Y$ with roots $-(a_{1}X_{1}+\cdots+a_{k}X_{k})$, $a_{1},\dots,a_{k}\in\mathbb{F}_{q}$; $(-1)^{k}\Delta(X_{1},\dots,X_{k})$ is the leading coefficient.
(ii) Differentiating both sides of (4.7) with respect to $Y$ gives
@@
For $11\leq s\leq q-$ 1, let $g_{s}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ be defined by
@@
(If $q^{n}>s+1$, then $g_{s}(x)=x^{q^{n}-1-s}$ for all $x\in\mathbb{F}_{q^{n}}$.) We remind the reader that $g_{s}$ here is different from the polynomial $g_{i,q}$ in (4.1). By (4.8),
@@
$\Delta=\sum_{i=0}^{k}(-1)^{i}Y^{q^{i}}\Delta_{i}=Y\Delta_{0}\Bigl{(}1-Y^{q-1}\frac{\Delta_{1}}{\Delta_{0}}+\text{terms of higer degree in $Y$}\Bigr{)}.$
@@
$\sum_{0\neq(a_{1},\dots,a_{k})\in\mathbb{F}_{q}^{k}}\frac{1}{(a_{1}X_{1}+\cdots+a_{k}X_{k})^{s}}=\begin{cases}-\displaystyle\frac{\Delta_{1}}{\Delta_{0}}&\text{if}\ s=q-1,\vspace{0.2em}\cr 0&\text{if}\ 1\leq s<q-1.\end{cases}$
@@
$\sum_{0\neq u\in E}\frac{1}{u^{s}}=\begin{cases}-\displaystyle\frac{\Delta_{1}(v_{1},\dots,v_{k})}{\Delta(v_{1},\dots,v_{k})^{q}}&\text{if}\ s=q-1,\vspace{0.2em}\cr 0&\text{if}\ 1\leq s<q-1.\end{cases}$
When $s<q-1$, $g_{s}$ is never $k$th order sum-free. Therefore, the “right” $q$-ary generalization of $f_{\text{\rm inv}}$ is $g_{q-1}$. From the above, we know that $g_{q-1}$ is not $k$th order sum-free if and only if there exist $v_{1},\dots,v_{k}\in\mathbb{F}_{q^{n}}$ such that $\Delta(v_{1},\dots,v_{k})\neq 0$ but $\Delta_{1}(v_{1},\dots,v_{k})=0$. We state this fact as a theorem since it is the foundation of some subsequent results.
@@
$g_{q-1}$ is not $k$th order sum-free if and only if there exist $v_{1},\dots,v_{k}\in\mathbb{F}_{q^{n}}$ such that $\Delta(v_{1},\dots,v_{k})\neq 0$ but $\Delta_{1}(v_{1},\dots,v_{k})=0$.
@@ 4.5. Generalized results for $g_{q-1}$
    Let $n\geq 2$. Then $g_{q-1}$ is $2$nd order sum-free if and only if $n$ is odd.
@@
In the above, $\Delta_{0}(v_{1},\dots,v_{k})=\Delta(v_{1},\dots,v_{k})^{q}$, and by (4.7),
@@
Theorem 4.10 ($q$-ary version of [5, Corollary 4]).
@@
In particular, $g_{q-1}$ is not $k$th order sum-free if and only if it is not $(n-k)$th order sum-free.
@@
If $11\leq k,l\leq n-$ 1 are such that $kl<n$ and $g_{q-1}$ on $\mathbb{F}_{q^{n}}$ is not $k$th order nor $l$th order sum-free, then $g_{q-1}$ is not $(k+l)$th order sum-free.
@@
(i) $\Rightarrow$ (ii). Since $x,\sigma^{2}(x),\dots,\sigma^{k}(x)$ are linearly dependent over $\mathbb{F}_{q}$, there exists $0\neq f=a_{k}X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}\in\mathbb{F}_{q}[X]$ such that $(f(\sigma))(x)=0$. We claim that $a_{k}\neq 0$ and $f\mid X^{n}-1$. Otherwise, $f_{1}:=\text{gcd}(f,X^{n}-1)$ has degree $<k$ and $(f_{1}(\sigma))(x)=0$. Then $x,\sigma(x),\dots,\sigma^{k-1}(x)$ are linearly dependent over $\mathbb{F}_{q}$, which is a contradiction. ∎
@@
If $X^{n}-1$ has a factor $X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}\in\mathbb{F}_{q}[X]$, then $g_{q-1}$ is not $k$th order sum-free.
@@ 4.6. Numerical results
$\Theta_{4}(X_{1},\dots,X_{4})$ is irreducible over $\overline{\mathbb{F}}_{2}$.
@@
$G(s_{1},\dots,s_{4})$ is irreducible in $\overline{\mathbb{F}}_{2}[s_{1},\dots,s_{4}]$.
@@
Clearly, $s_{1}^{4}+s_{2}^{2}+s_{1}s_{3}$ is irreducible in $\overline{\mathbb{F}}_{2}[s_{1},s_{2},s_{3}]$. Hence $A=a(s_{1}^{4}+s_{2}^{2}+s_{1}s_{3})^{i}$ and $B=b(s_{1}^{4}+s_{2}^{2}+s_{1}s_{3})^{2-i}$ for some $a,b\in\overline{\mathbb{F}}_{2}^{*}$ and $i\in\{0,,1,2\}$. In any case, $A+B$ does not contain any monomial of degree $3$. Thus $A+B\neq s_{1}(s_{1}+s_{3})$, which is a contradiction. ∎
@@
$H(X_{1},X_{2},X_{3})$ is irreducible over $\overline{\mathbb{F}}_{2}$.
@@
Assume to the contrary that $\Theta_{4}$ has an absolutely irreducible factor $f$ of degree $d$ with $80<d<$ 8. Treating $f$ as a polynomial in $X_{4}$ over $\overline{\mathbb{F}}_{2}[X_{1},X_{2},X_{3}]$, we may write
@@
where $A_{i}\in\overline{\mathbb{F}}_{2}[X_{1},X_{2},X_{3}]$ is homogeneous of degree $d-i$. Then $A_{0}\mid\Theta_{4}(X_{1},X_{2},X_{3},0)=H^{2}$. Since $H$ is absolutely irreducible (Lemma A3) and $\deg H=4$, we must have $d=4$ and $A_{0}=cH$ for some $c\in\overline{\mathbb{F}}_{2}$, so

### v2

@@ 4. Generalizations
Recall that a function $f:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is said to be $k$th order sum-free if $\sum_{x\in A}f(x)\neq 0$ for all $k$-dimensional $\mathbb{F}_{q}$-affine subspaces $A$ of $\mathbb{F}_{q^{n}}$. In general, compared with their binary counterpart, $q$-ary sum-free functions are more difficult to characterize, construct and analyze. APN functions in odd characteristic, studied in [9, 10], do not represent the 2nd order $q$-ary sum-free functions. On the other hand, the generalized almost perfect nonlinear (GAPN) functions in odd characteristic, introduced and explored in [15], appear to be related to the 1st order sum-free functions. However, this connection will not be pursued in the present paper.
We first take a look of the 1st order and the (n1)(n-1)th order $q$-ary sum-free functions.
@@ 4.1. The 1st order
Let $f(X)=\sum_{i=0}^{q^{n}-1}a_{i}X^{i}\in\mathbb{F}_{q^{n}}[X]$. Then $f:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is 11st order sum-free if and only if
@@
Example 4.2.
Let $q=3$ and $n=2$. We have
$g_{i,3}(X)=\begin{cases}0&\text{if}\ i=0,1,3,\cr-1&\text{if}\ i=2,4,6,\cr-X&\text{if}\ i=5,\cr X&\text{if}\ i=7,\cr-X^{2}-1&\text{if}\ i=8.\end{cases}$
For $0\leq i\leq 8$, $\gcd(g_{i,3}(X^{3}-X),X^{9}-X)=1$ if and only if $i=2,4,6$. Hence $X^{i}$ is 1st order sum-free on $\mathbb{F}_{3^{2}}$ if and only if $i\in\{2,4,6\}$.
4.2. The (n1)(n-1)th order
@@
$\kern 85.35826pt\cdot\Bigl[\sum_{i=0}^{q^{n}-1}a_{i}\sum_{x\in\mathbb{F}_{q^{n}}}x^{i+j_{0}q^{0}+\cdots+j_{n-1}q^{n-1}}\Bigr]$
@@
In the above notation, $f$ is (n1)(n-1)th order sum-free if and only if
@@
Condition (4.5) can always be satisfied since the right side of inequality (4.5) takes at most $q^{n}-1$ values as $b$ runs over $\mathbb{F}_{q^{n}}^{*}$. Condition (4.6) is difficult to verify in general. However, in special cases, condition (4.6) can be made more explicit.
Let $0<t\leq q-1$ and assume that for $0<j_{0}+\cdots+j_{n-1}\leq q-1$,
@@
[…]
@@
$(j_{0},\dots,j_{n-1})=(0,\dots,0,\overset{i}{t},0\,\dots,0)\quad\text{for some }\ 0\leq i\leq n-1.$
Then the left side of inequality (4.6) becomes
@@
is a $q$-polynomial. Therefore, if $L_{t}(X)$ is a permutation polynomial of $\mathbb{F}_{q^{n}}$, then condition (4.6) is satisfied.
Example 4.4.
Let $q=5$, $n=2$, and $\mathbb{F}_{5^{2}}=\mathbb{F}_{5}(\alpha)$, where $\alpha^{2}=2$. Choose a permutation $q$-polynomial of $\mathbb{F}_{q^{2}}$, say $X+\alpha X^{5}$. Let $t=3$ and choose $a_{24-3}$ and $a_{24-3\cdot 5}$ such that
$X+\alpha X^{5}=L_{3}(X)=a_{24-3}X+a_{24-3\cdot 5}X^{5},$
i.e., $a_{21}=1$ and $a_{9}=\alpha$. For $i\in\{0,\dots,23\}\setminus\{9,21\}$, set $a_{i}=0$. Then codition (4.6) is satisfied. Moreover, the right side of (4.5) is 0. Therefore, for all $a_{24}\in\mathbb{F}_{5^{2}}^{*}$,
$f(X)=\alpha X^{9}+X^{21}+a_{24}X^{24}$
is 1st order sum-free on $\mathbb{F}_{5^{2}}$.
@@
4.3. Formulas about the Moore determinant
Recall that the $q$-ary Moore determinant $\Delta(X_{1},\dots,X_{k})$ and its relatives $\Delta_{i}(X_{1},$ $\dots,X_{k})$ are defined in §2.3.
@@ 4.3. Formulas about the Moore determinant
(i) $\Delta(Y,X_{1},\dots,X_{k})$ is a polynomial of degree $q^{k}$ in $Y$ with roots $(a_{1}X_{1}+\cdots+a_{k}X_{k})$, $a_{1},\dots,a_{k}\in\mathbb{F}_{q}$; $(-1)^{k}\Delta(X_{1},\dots,X_{k})$ is the leading coefficient.
(ii) Differentiating both sides of equation (4.7) with respect to $Y$ gives
@@
For $1\leq s\leq q-1$, let $g_{s}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ be defined by
@@
(If $q^{n}>s+1$, then $g_{s}(x)=x^{q^{n}-1-s}$ for all $x\in\mathbb{F}_{q^{n}}$.) We remind the reader that $g_{s}$ here is different from the polynomial $g_{i,q}$ in equation (4.1). By equation (4.8),
@@
$\Delta=\sum_{i=0}^{k}(-1)^{i}Y^{q^{i}}\Delta_{i}=Y\Delta_{0}\Bigl(1-Y^{q-1}\frac{\Delta_{1}}{\Delta_{0}}+\text{terms of higher degree in $Y$}\Bigr).$
@@
$\sum_{0\neq(a_{1},\dots,a_{k})\in\mathbb{F}_{q}^{k}}\frac{1}{(a_{1}X_{1}+\cdots+a_{k}X_{k})^{s}}=\begin{cases}-\displaystyle\frac{\Delta_{1}}{\Delta_{0}}&\text{if}\ s=q-1,\vskip 1.99997pt\cr 0&\text{if}\ 1\leq s<q-1.\end{cases}$
@@
$\sum_{0\neq u\in E}\frac{1}{u^{s}}=\begin{cases}-\displaystyle\frac{\Delta_{1}(v_{1},\dots,v_{k})}{\Delta(v_{1},\dots,v_{k})^{q}}&\text{if}\ s=q-1,\vskip 1.99997pt\cr 0&\text{if}\ 1\leq s<q-1.\end{cases}$
When s<q1s<q-1, $g_{s}$ is never $k$th order sum-free. Therefore, the “right” $q$-ary generalization of $f_{\text{\rm inv}}$ is $g_{q-1}$. From the above, we know that $g_{q-1}$ is not $k$th order sum-free if and only if there exist $v_{1},\dots,v_{k}\in\mathbb{F}_{q^{n}}$ such that $\Delta(v_{1},\dots,v_{k})\neq 0$ but $\Delta_{1}(v_{1},\dots,v_{k})=0$. We state this fact as a theorem since it is the foundation of some subsequent results.
@@
The function $g_{q-1}$ is not $k$th order sum-free if and only if there exist $v_{1},\dots,v_{k}\in\mathbb{F}_{q^{n}}$ such that $\Delta(v_{1},\dots,v_{k})\neq 0$ but $\Delta_{1}(v_{1},\dots,v_{k})=0$.
@@ 4.5. Generalized results for $g_{q-1}$
    Let $n\geq 2$. Then $g_{q-1}$ is 22nd order sum-free if and only if $n$ is odd.
@@
For $n\geq 2$, define
When $q=2$, $\mathcal{K}_{n,2}=\mathcal{K}_{n}$, which is defined in equation (1.1).
@@
In the above, $\Delta_{0}(v_{1},\dots,v_{k})=\Delta(v_{1},\dots,v_{k})^{q}$, and by equation (4.7),
@@
Theorem 4.10 ($q$-ary version of [5, Theorem 3]).
@@
In particular, $k\in\mathcal{K}_{n,q}$ if and only if $n-k\in\mathcal{K}_{n,q}$.
@@
If $k,l\in\mathcal{K}_{n,q}$ and $kl<n$, then $k+l\in\mathcal{K}_{n,q}$.
@@
(i) $\Rightarrow$ (ii). Since $x,\sigma^{2}(x),\dots,\sigma^{k}(x)$ are linearly dependent over $\mathbb{F}_{q}$, there exists $0\neq f=a_{k}X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}\in\mathbb{F}_{q}[X]$ such that $(f(\sigma))(x)=0$. We claim that $a_{k}\neq 0$ and $f\mid X^{n}-1$. Otherwise, $f_{1}:=\text{gcd}(f,X^{n}-1)$ has degree <k<k and $(f_{1}(\sigma))(x)=0$. Then $x,\sigma(x),\dots,\sigma^{k-1}(x)$ are linearly dependent over $\mathbb{F}_{q}$, which is a contradiction. ∎
@@
If $X^{n}-1$ has a factor $X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}\in\mathbb{F}_{q}[X]$, then $k\in\mathcal{K}_{n,q}$.
@@ 4.6. Numerical results
The polynomial $\Theta_{4}(X_{1},\dots,X_{4})$ is irreducible over $\overline{\mathbb{F}}_{2}$.
@@
The polynomial $G(s_{1},\dots,s_{4})$ is irreducible in $\overline{\mathbb{F}}_{2}[s_{1},\dots,s_{4}]$.
@@
Clearly, $s_{1}^{4}+s_{2}^{2}+s_{1}s_{3}$ is irreducible in $\overline{\mathbb{F}}_{2}[s_{1},s_{2},s_{3}]$. Hence $A=a(s_{1}^{4}+s_{2}^{2}+s_{1}s_{3})^{i}$ and $B=b(s_{1}^{4}+s_{2}^{2}+s_{1}s_{3})^{2-i}$ for some $a,b\in\overline{\mathbb{F}}_{2}^{*}$ and $i\in\{0,1,2\}$. In any case, $A+B$ does not contain any monomial of degree $3$. Thus $A+B\neq s_{1}(s_{1}s_{2}+s_{3})$, which is a contradiction. ∎
@@
The polynomial $H(X_{1},X_{2},X_{3})$ is irreducible over $\overline{\mathbb{F}}_{2}$.
@@
Assume to the contrary that $\Theta_{4}$ has an absolutely irreducible factor $f$ of degree $d$ with $0<d<8$. Treating $f$ as a polynomial in $X_{4}$ over $\overline{\mathbb{F}}_{2}[X_{1},X_{2},X_{3}]$, we may write
@@
where $A_{i}\in\overline{\mathbb{F}}_{2}[X_{1},X_{2},X_{3}]$ is homogeneous of degree did-i. Then $A_{0}\mid\Theta_{4}(X_{1},X_{2},X_{3},0)=H^{2}$. Since $H$ is absolutely irreducible (Lemma A3) and $\deg H=4$, we must have $d=4$ and $A_{0}=cH$ for some $c\in\overline{\mathbb{F}}_{2}$, so